# Enhanced Optimization Strategies to Design an Underactuated Hand Exoskeleton

### v1

@@ Abstract
Exoskeletons can boost human strength and provide assistance to individuals with physical disabilities. However, ensuring safety and optimal performance in their design poses substantial challenges. This study presents the design process for an underactuated hand exoskeleton (U-HEx), first including a single objective (maximizing force transmission), then expanding into multi objective (also minimizing torque variance and actuator displacement). The optimization relies on a Genetic Algorithm, the Big Bang-Big Crunch Algorithm, and their versions for multi-objective optimization. Analyses revealed that using Big Bang-Big Crunch provides high and more consistent results in terms of optimality with lower convergence time. In addition, adding more objectives offers a variety of trade-off solutions to the designers, who might later set priorities for the objectives without repeating the process – at the cost of complicating the optimization algorithm and computational burden. These findings underline the importance of performing proper optimization while designing exoskeletons, as well as providing a significant improvement to this specific robotic design.
Index Terms:
@@
Prosthetics and Exoskeletons, Rehabilitation Robotics, Mechanism Design, Evolutionary Computation and Optimization.
I Introduction
@@
Exoskeletons are wearable robotic devices that can improve users’ functional abilities during physically challenging tasks [1, 2, 3] or assist/rehabilitate them in case of functional disabilities [4, 5, 6]. They are usually designed to be worn on a specific body location; thus, their proper design is crucial for force transmission efficiency but also, and most importantly, for users’ safety [7]. For an exoskeleton to be safe, designers must consider the alignment between the anatomic and mechanical joints and possible (intentional or unintentional) movements of users during operation [8, 9]. In addition, the exoskeleton design must be optimized to provide the highest level of well-distributed output forces through small and lightweight actuators.
@@
The engineering design process is usually performed through mathematical optimization, which is the search for the best element (optimum) within a set of alternatives (decision variables’ values) based on one or more specific criteria (objectives) [10, 11, 12]. Traditional optimization methods rely on numerical and calculus-based techniques [13] and might be challenging for solving exoskeleton design [14] due to properties such as non-discrete domains, non-differentiability, multimodality, discontinuity [15], and the need for reliability [16] and robustness [17]. Alternatively, the nature-inspired Evolutionary Computation (EC) methods – particularly Evolutionary Algorithms (EAs) – are well-known and effective in dealing with engineering optimization problems [18] and often with exoskeleton design [14].
@@
Previously, we proposed optimizing an underactuated hand exoskeleton (U-HEx [19], shown in Fig. 1) using EAs, in which we retrieved the best set of link lengths (decision variables) that maximize the output force transmission (objective function) [20]. While the first design of U-HEx was optimized with a naive-deterministic search approach, which iteratively analyzed each possible combination of design parameters (i.e., “brute force”), we employed a Genetic Algorithm (GA) [21, 22] and the Big Bang-Big Crunch algorithm (BBBC) [23]. Comparative analyses revealed that both EAs yielded more consistent and optimal solutions than brute force in a significantly shorter time. This allowed us to extend the search space by including more decision variables to the problem (i.e., three more link lengths in the design), significantly improving U-HEx’s efficiency.
However, our previous work had a few limitations. Firstly, we only investigated the impact of different optimization methods over a single objective function. Secondly, although we included mechanical constraints to ensure the device’s safety and operational efficiency, we did not constrain the desired actuator movements to reach a natural range of motion for the user’s hand. In other words, we have found an optimal set of link lengths with no prior assumption on the desired actuator displacements. This assumption cannot be overlooked, since our design involves a linear actuator with a physical limit to how much it could move from its base. In fact, choosing the suitable actuator for an exoskeleton design is a challenging step involving many factors [24]. For the U-HEx design, a linear DC motor was selected due to its low cost and compact design – so that it can fit on top of the hand [19]. Therefore, it is crucial to minimize the desired range of actuator displacement to assist natural ranges of motion for the finger joints without sacrificing the force transmission performance. On the other hand, we have previously constrained the ratio between the actuated finger joints in our optimization problem, while in reality, it would be preferred to be as close to 1 for well-distributed and safe force transmission.
In this study, we have built up on our previous single-objective optimization problem (SOOP) [20] by improving U-HEx’s design in two alternative design processes:
@@
    we redefined the SOOP with an additional constraint to the maximum linear actuator displacement in parallel to the actuator chosen for the original design [19], and we hypothesize that this would result in a lower overall transmission and potentially different trends between different optimization methods – even though we have very effective optimization methods in hand ($H_{1}$); and
@@
    we extended the SOOP to a multi-objective optimization problem (MOOP) by also optimizing the balance between the torques applied to each finger joint and minimizing the linear displacement needed by the actuator (i.e., turning constraints into objectives), and we hypothesize that treating constraints as further objectives would better accommodate the problem and lead to a trade-off set of different designs balancing the device’s properties ($H_{2}$).
@@
Our contributions involve (i) providing a set of optimized designs for a state-of-the-art exoskeleton that was originally optimized with inefficient methods [19], (ii) providing a walkthrough for a robotic design process that can be used as a guideline by future and less experienced robotic designers, (iii) comparing different algorithms and strategies for MOOPs to determine which one is most suitable for our design problem, and (iv) introducing two novel EC methods (NS-BBBC and SP-BBBC) suitable for robotic optimization.
@@
The rest of the work is organized as follows: Sec. III describes and discusses the first design process, which optimizes U-HEx only considering a single objective (i.e., maximize the force transmission); Sec. IV describes and discusses the second design process, which optimizes U-HEx by adding two more objectives (i.e., distribute the distribution of forces to the finger and minimize the linear actuator displacement); Sec. V discusses the differences in the retrieved designs and the impact of each optimization process; and Sec. VI summarizes the work and presents future ideas.
Figure 2: Schematic of the optimization process for U-HEx: (a) its optimization problem; (b) its kinematics and statics to analyze its operation in terms of movements and force transmission. All the optimization algorithms are implemented in MATLAB. For a given set of link lengths, a Simulink model is executed to compute numerical inverse kinematics and analytical statics as the finger joints are iterated from fully open to fully closed.
II Exoskeleton Design Optimization
II-A U-HEx Kinematics
@@
U-HEx is a hand exoskeleton designed for hand rehabilitation, whose kinematic structure was detailed in its original work [19] and depicted in Fig. 2(a). Each gray dot represents a passive rotational joint (or anatomical finger joint) while empty circles ($H$, $E$, and $C$) are fixed points along rigid links with $90^{o}$ fixed angle. The system has only one linear actuator between the points $O$ and $A$. The exoskeleton is fixed on the hand from points $K$ and $O$ with variable lengths in the $x$ and $y$ directions. Finally, the exoskeleton is attached to the finger phalanges from points $I$ and $J$ with passive linear joint sliders represented as $c_{1}$ and $c_{2}$, respectively.
II-B Optimization Schema
@@
U-HEx design is optimized to retrieve the optimal link lengths following the schema and design process summarized in Fig. 2. The decision variables of the problem are the link lengths (depicted as red lines in the figure). Specifically, the original design of U-HEx [19] was optimized considering only six link lengths: $\overline{BC}$, $\overline{CD}$, $\overline{DE}$, $\overline{EF}$, $\overline{FG}$, and $\overline{GH}$. This limitation was caused by the heavy computational requirements of brute force, which made it impossible to run the optimization with more link lengths. In our previous work [20], as well as in the current work, we included an additional three link lengths ($\overline{BK}$, $\overline{CI}$, and $\overline{EJ}$), potentially creating a more optimal solution. The bounds of the decision variables (i.e., the range of the link lengths) are reported in Table I and were chosen empirically in U-HEx’s original work [19].
TABLE I: Link Length Ranges / Decision Variable Bounds
@@
At each step of the process, a new set of link lengths is assigned by the optimization method in MATLAB. As shown in Fig. 2 (b), this set is sent to a Simulink model (with a fixed-step solver) to compute the numerical inverse kinematics [19] as the finger joints are iterated from fully open ($0^{o}$) to fully closed ($80^{o}$ for MCP and $90^{o}$ for PIP). This process results in the corresponding movements around the passive joints (including the linear sliders $c_{1}$, $c_{2}$) and the desired actuator displacement ($L_{x}$). These movements are then used for the statics equations through analytical methods to compute the force transmission – i.e., expected torques around the finger joints ($\tau_{MCP},\tau_{PIP}$) for a unit actuator force (1 N).
Obtaining the force transmission results terminates the Simulink file, while the MATLAB file evaluates the obtained performance in terms of the constraints and the optimality. Depending on the performance, the optimization algorithm repeats this loop until the requirements are satisfied.
II-C Optimization Algorithms (Evolutionary Algorithms)
@@
In this paper, we will explore two optimization problems (SOOP and MOOP) using the same optimization algorithms based on EC. EC is a sub-field of soft computing offering nature-inspired algorithms: instead of updating on a single point in the search space, it generates a population of potential solutions (i.e., a set of different combinations of decision variables’ values) and evolves them toward the optimum using different metaheuristics (e.g., inspired by genetic recombination, natural selection, universe creation, animal behavior when foraging for food). Since the generic framework of EC methods does not rely on a specific mathematical model and can also tackle objective functions defined implicitly through simulations, they are particularly suitable for engineering problems such as optimizing the design of U-HEx. Specifically, we will employ two different EAs, as described below.
II-C1 Genetic Algorithm (GA)
@@
GAs are the most widely used EAs because they directly replicate the process of natural selection and the concept of survival of the fittest [25, 22]. As summarized in Algorithm 1, GA (i) creates a population of random solutions $P$ within the problem’s search space, (ii) assigns a fitness value to each solution based on the objective function, and (iii) produces new solutions $Q$ by combining the values of the solutions in the current population through a process known as crossover. Only the fittest solutions $M$ undergo crossover and are retained in subsequent generations, which allows the GA to evolve its population towards the problem’s optimal solution. Similar to biological processes, the crossover operation utilizes the characteristics of high-quality solutions (parents) to generate new, comparable solutions (offspring), thereby accelerating convergence to an optimum. However, there is no guarantee that the retrieved optimum is global rather than local. To address this issue, GAs incorporate an additional operator inspired by genetic mutation. This randomly alters the values of newly generated solutions, enhances search-space exploration, and helps in avoiding local optima.
@@
output : The most fitting solution $\textnormal{{P}}(1)$
@@
II-C2 Big Bang-Big Crunch Algorithm (BBBC)
BBBC [23] is inspired by the universe’s evolution through two phases: explosion and implosion. The process involves energy dissipation, which creates disorder and randomness, followed by the reorganization of this randomness into a new order. As summarized in Algorithm 2, BBBC (i) generates an initial random population $P$ uniformly distributed across the search space (the explosion, or “big bang”), and (ii) evaluates and collects the population into their center of mass $cm$ (the implosion, or “big crunch”). As the iterations progress, these two phases are repeated, with new solutions being generated closer to the center of mass, converging to the problem’s optimal solution. BBBC is recognized for its superior convergence speed compared to GAs, making it a suitable choice for our problem, given the high runtime of the objective function caused by the execution of a Simulink simulation.

### v2

@@ Abstract
Exoskeletons can boost human strength and provide assistance to individuals with physical disabilities. However, ensuring safety and optimal performance in their design poses substantial challenges. This study presents the design process for an underactuated hand exoskeleton, first including a single objective (maximizing force transmission), then expanding into multi-objective (also minimizing torque variance and actuator displacement). The optimization relies on a Genetic Algorithm, the Big Bang-Big Crunch Algorithm, and their versions for multi-objective optimization. Analyses revealed that using Big Bang-Big Crunch provides high and more consistent results in terms of optimality with lower convergence time. In addition, adding more objectives offers a variety of trade-off solutions to the designers, who might later set priorities for the objectives without repeating the process – at the cost of complicating the optimization algorithm and computational burden. These findings underline the importance of performing proper optimization while designing exoskeletons, as well as providing a significant improvement to this specific robotic design.
keywords:
@@
Hand Exoskeleton , Rehabilitation Robotics , Evolutionary Computation , Multi-Objective Optimization , Design Optimization
journal: ArXiv Preprint
[…]
1 Introduction
@@
Exoskeletons are wearable robotic devices that can improve users’ functional abilities during physically challenging tasks [poliero2020applicability, butzer2021fully] or assist/rehabilitate them in case of functional disabilities [stroppa2017robot, huang2015design, pons2010rehabilitation]. They are usually designed to be worn on a specific body location; thus, their proper design is crucial for force transmission efficiency but also, and most importantly, for users’ safety [sarac2019design]. For an exoskeleton to be safe, the anatomic and mechanical joints must be aligned during possible (intentional or unintentional) movements of users [van2015considerations, fisahn2016effectiveness]. In addition, the exoskeleton design must be optimized to provide the highest level of well-distributed output forces through small and lightweight actuators.
@@
The engineering design is usually performed through mathematical optimization, searching for the best element (optimum) within a set of alternatives (decision variables’ values) based on one or more specific criteria (objectives) [sioshansi2017optimization, statnikov2012multicriteria, andersson2000survey]. Traditional optimization methods rely on numerical and calculus-based techniques [bonnans2006numerical] and might be challenging for exoskeleton design [stroppa2023optimizing] due to properties such as non-discrete domains, multimodality, discontinuity [norde2000characterizing], and the need for reliability [dizangian2015reliability] and robustness [nomaguchi2016robust]. Alternatively, the nature-inspired Evolutionary Computation (EC) methods – particularly Evolutionary Algorithms (EAs) – are well-known and effective in dealing with engineering optimization problems [dumitrescu2000evolutionary, wang2017autonomous] and often with exoskeleton design [stroppa2023optimizing, datta2015analysis] and other robotic devices [stroppa2024optimizing].
@@
Previously, we proposed optimizing an underactuated hand exoskeleton (U-HEx [sarac2017design], shown in Fig. 1) using EAs, in which we retrieved the best set of link lengths (decision variables) that maximize the output force transmission (objective function) [akbas2024impact]. We specifically chose EAs over other nonlinear numerical methods due to their robustness in handling black-box optimization and their ability to efficiently explore the design space of our underactuated hand exoskeleton without requiring explicit gradients or analytical formulations for the objective functions – especially considering that the objective functions come from a motion simulation of the kinematic chain. While the first design of U-HEx was optimized with a naive-deterministic search approach, which iteratively analyzed each possible combination of design parameters (i.e., “brute force”), we employed a Genetic Algorithm (GA) [goldberg1987genetic, goldberg1990real] and the Big Bang-Big Crunch algorithm (BBBC) [erol2006new]. Comparative analyses revealed that both EAs yielded more consistent and optimal solutions than brute force in a significantly shorter time. This allowed us to extend the search space by including more decision variables (i.e., three more link lengths), significantly improving U-HEx’s efficiency.
Yet, our previous work had a few limitations. Firstly, we only focused on a single objective function. Secondly, we did not constrain the desired actuator movements to reach a natural range of motion for the user’s hand. In other words, we have found an optimal set of link lengths with no prior assumption on the desired actuator displacements. This assumption cannot be overlooked, since our design involves a linear actuator with a physical limit to how much it could move from its base – so that it can fit on top of the hand [sarac2017design]. Therefore, it is crucial to minimize the desired range of actuator displacement to assist natural ranges of motion for the finger joints without sacrificing the force transmission performance. On the other hand, we have previously constrained the ratio between the actuated finger joints in our optimization problem, while it would be preferred to be as close to 1 for well-distributed and safe force transmission.
In this study, we have built up on our Single-Objective Optimization Problem (SOOP) [akbas2024impact] by improving U-HEx’s design in two alternative design processes:
@@
    we redefined the SOOP with an additional constraint to the actuator displacement based on the original design [sarac2017design], and we hypothesize that this would result in a lower overall transmission and potentially different trends between optimization methods – even though we have very effective optimization methods in hand ($H_{1}$); and
@@
    we extended the SOOP to a Multi-Objective Optimization Problem (MOOP) by also optimizing the balance between the torques applied to each finger joint and minimizing the required actuator displacement (i.e., turning constraints into objectives), and we hypothesize that treating constraints as further objectives would better accommodate the problem and lead to a trade-off set of different designs balancing the device’s properties ($H_{2}$).
@@
Our contributions involve (i) providing a set of optimized designs for a state-of-the-art exoskeleton that was originally optimized with inefficient methods [sarac2017design], (ii) providing a walkthrough for a robotic design process that can be used as a guideline by future and less experienced robotic designers, (iii) comparing different algorithms and strategies for MOOPs to determine which one is most suitable for our design problem, and (iv) introducing two novel EC methods (NS-BBBC and SP-BBBC) suitable for robotic optimization, which are based on a state-of-the-art algorithm for single-objective optimization (i.e., BBBC). Table 1 summarizes the different contributions on each work of the U-HEx device, comparing the current work with its state-of-the-art.
Table 1: Comparison of different works novelties on U-HEx
Reference Novelty Optimization Method
Sarac et al. 2017 [sarac2017design] Kinematics and mechanical design Brute Force
Akbas et al. 2024 [akbas2024impact] Optimized device based on maximum torque magnitude Brute Force, GA, BBBC
Current Work Comprehensive and broader optimization of the device on multiple criteria GA, BBBC, NSGA-II, SPEA2, NS-BBBC, SP-BBBC
@@
The rest of the work is organized as follows: Sec. 2 defines the optimization design and provides a background on the methods we used; Sec. 3 describes the first optimization problem with a single objective (i.e., maximize the force transmission); Sec. 4 describes the second optimization problem with two more objectives (i.e., minimize the distribution of forces and actuator displacement); Sec. LABEL:sec:comparison discusses the differences in the retrieved designs and the impact of each optimization process; and Sec. LABEL:sec:conclusion summarizes the work and presents future ideas.
2 Exoskeleton Design Optimization
2.1 U-HEx Kinematics
U-HEx in Fig. 1 is a hand exoskeleton designed for rehabilitation [sarac2017design]. It was designed as a linkage structure with mechanical joints intentionally misaligned from the anatomical joints of the user. This design choice has multiple advantages. Firstly, the misalignment offers a much safer operation, since the mechanical device does not impose hard forces on the joints directly. Any sort of undesired or extreme force would be transmitted to the fingers in a more natural way without the risk of harming the user. Secondly, thanks to this misalignment, U-HEx can adapt its operation automatically to users with different finger sizes as well. Finally, since U-HEx joints do not need to be aligned with the anatomical joints strictly, it is noticeably easy to wear.
The misalignments of U-HEx were designed carefully to achieve underactuation. A linear DC motor allows the device to close/open the finger in a direct manner. The idea behind underactuation is simple: the actuation is turned into motion at the first output joint until the first mechanical link gets in contact with an object. After the first link reaches an object and starts feeling the interaction forces, the spring-like joint coupling between output joints allows the mechanism to transmit the same actuation for the second output joint. Actuator forces are transmitted to output joints until the grasping is completed [Denizon2024, Yan2022]. From the point of hand exoskeletons, users’ fingers as a part of the mechanism: anatomical coupling between finger joints defines the required spring-like behavior for the underactuation [Li2023, Gu2024, sarac2017design]. Then, efficient force transmission requires the use of additional passive joints, which can assist motion for finger joints without aligning mechanical joints to the finger joints.
Fig. 2 represents the operation flow of U-HEx. The finger component is connected to the user’s finger and assists the finger in flexing until both finger joints are fully extended and the actuator stroke is at its minimum limit. In the beginning, as the actuator starts to move, the exoskeleton transmits forces to the proximal phalange and rotates the MCP joint. Encountering external forces along the proximal phalange caused by the physical interaction finalizes the rotation along the MCP joint. These forces transmit actuation forces to the intermediate phalange through the linkage-based mechanism with passive joints and rotating the PIP joint. Having two finger joints to control, the grasping task is completed when the PIP joint is rotated enough for the intermediate phalange to reach the object. Similarly, the extension of the finger starts by rotating the PIP joint first and the MCP joint later until the finger is totally extended and finger joints reach their physical limits. As a safety measure, the displacement stroke is designed to reach its minimum limit when both finger joints are extended to their natural limits.
Figure 2: Underactuation concept of U-HEx during a grasping task: (left) initial pose with all finger joints extended, (mid) actuator force moves the MCP joint until the first finger phalange gets in touch with the object, (right) actuator force is transmitted to move the PIP joint when the first finger phalange touches the object and finally grasping task is completed when both phalanges touch the object
@@ 2.1 U-HEx Kinematics
Fig. 3(a) represents the kinematic structure of U-HEx. Each gray dot represents a passive rotational joint (or anatomical joint) while empty circles ($H$, $E$, and $C$) are fixed points along rigid links with $90^{o}$ fixed angle. The system has only one linear actuator between the points $O$ and $A$. The exoskeleton is fixed on the hand from points $K$ and $O$ with variable lengths in the $x$ and $y$ directions. Finally, the exoskeleton is attached to the finger phalanges from points $I$ and $J$ with passive linear joint sliders represented as $c_{1}$ and $c_{2}$, respectively.
Figure 3: Optimization process for U-HEx: for a given set of link lengths, a Simulink model is executed to compute numerical inverse kinematics and analytical statics as the finger joints are iterated from fully open to fully closed; while the optimization algorithms are implemented in MATLAB.
@@
U-HEx design is optimized to retrieve the optimal link lengths following the schema and design process summarized in Fig. 3. The decision variables of the problem are the link lengths (depicted as red lines in the figure). The original U-HEx [sarac2017design] was optimized only for six link lengths: $\overline{BC}$, $\overline{CD}$, $\overline{DE}$, $\overline{EF}$, $\overline{FG}$, and $\overline{GH}$. due to the heavy computational requirements of brute force, which made it impossible to run the optimization with more link lengths. In our previous work [akbas2024impact], we included an additional three link lengths ($\overline{BK}$, $\overline{CI}$, and $\overline{EJ}$) in the optimization problem, potentially creating a more optimal solution. The bounds of the decision variables (i.e., the range of the link lengths) are reported in Table 2 and were chosen empirically [sarac2017design].
Table 2: Link Length Ranges / Decision Variable Bounds
@@
Fig. 3 shows that at each step of the process, a new set of link lengths is assigned by the optimization method in MATLAB. This set is used to compute the numerical inverse kinematics [sarac2017design] as the finger joints are iterated from fully open ($0^{o}$) to fully closed ($80^{o}$ MCP; $90^{o}$ PIP) with a fixed-step solver. This process results in the corresponding movements around the joints (e.g., linear sliders $c_{1}$, $c_{2}$) and the desired actuator displacement ($L_{x}$). These movements are then used to compute the force transmission – i.e., expected torques around the finger joints ($\tau_{MCP},\tau_{PIP}$) for a unit actuator force (1 N) through analytical statics equations. Obtaining the force transmission triggers the MATLAB file to evaluate the obtained performance in terms of the constraints and the optimality. Depending on the performance, the optimization algorithm repeats this loop until the requirements are satisfied.
2.2 Optimization Algorithms (Evolutionary Algorithms)
@@
In this paper, we will explore two optimization problems (SOOP and MOOP) using the same optimization algorithms based on EC. EC is a sub-field of soft computing offering nature-inspired algorithms: instead of updating on a single point in the search space, it generates a population of potential solutions (i.e., a set of different combinations of decision variables’ values) and evolves them toward the optimum using different metaheuristics (e.g., inspired by genetic recombination, natural selection, universe creation). Since the generic framework of EC methods does not rely on a specific mathematical model and can tackle objective functions defined implicitly through simulations, they are suitable for engineering problems such as optimizing U-HEx. Specifically, we will employ two different EAs, as described below.
2.2.1 Genetic Algorithm (GA)
@@
GAs are the most widely used EAs because they directly replicate the process of natural selection and the concept of survival of the fittest [goldberg1989genetic, goldberg1990real]. As summarized in Algorithm 1, GA (i) creates a population of random solutions $P$ within the search space, (ii) assigns a fitness value to each solution based on the objective function, and (iii) produces new solutions $Q$ by combining the values of the solutions in the current population through a crossover process. Only the fittest solutions $M$ undergo crossover and are retained in subsequent generations, which allows the GA to evolve its population towards the optimal solution. Similar to biological processes, the crossover operation utilizes the characteristics of high-quality solutions (parents) to generate new, comparable solutions (offspring) – accelerating convergence to an optimum. However, there is no guarantee that the retrieved optimum is global rather than local. To address this issue, GAs incorporate an additional operator inspired by genetic mutation. This randomly alters the values of newly generated solutions, enhances search-space exploration, and helps in avoiding local optima.
@@
output : The most fitting solution $\textnormal{{P}}\left(1\right)$
@@
2.2.2 Big Bang-Big Crunch Algorithm (BBBC)
BBBC [erol2006new] is inspired by the universe’s evolution through two phases: explosion and implosion. The process involves energy dissipation, which creates disorder and randomness and then reorganizes this randomness into a new order. As summarized in Algorithm 2, BBBC (i) generates an initial random population $P$ uniformly distributed across the search space (the explosion, or “big bang”), and (ii) evaluates and collects the population into their center of mass $cm$ (the implosion, or “big crunch”). As the iterations progress, these two phases are repeated, with new solutions being generated closer to the center of mass, converging to the optimal solution. BBBC is recognized for its superior convergence speed compared to GAs, making it a suitable choice for our problem, given the high runtime of the objective function executed in a Simulink simulation.